OpenStudy (anonymous):

4 x 3. will medal so pls help :-(

OpenStudy (anonymous):

Hey there!

OpenStudy (anonymous):

Multiplication is very similar to addition!

OpenStudy (anonymous):

Let me give you an example!! 2 x 5 = 5 + 5 = 10 But, you can also do it another way! 2 x 5 = 2 + 2 + 2 + 2 + 2 = 10

OpenStudy (anonymous):

Basically, 2 times 5 means adding 5 by itself two times, or adding 2 by itself 5 times!

OpenStudy (anonymous):

So to find 4 x 3, we can solve one of these additions: 4 x 3 = 3 + 3 + 3 + 3 = ?? 4 x 3 = 4 + 4 + 4 = ??

OpenStudy (anonymous):

Can you do that? @RozieDoze

OpenStudy (anonymous):

@RosieDoze *

OpenStudy (anonymous):

12?????

OpenStudy (anonymous):

Correct! :)

OpenStudy (anonymous):

Great job!!